Agenda/Minutes of the VIS Meeting on 2020/5/29
2020/5/29 14:00 -
Participants: Takahashi, Fabian, Sato, Hirata, Shoda, Miyo, Tanaka, Aso, Lucia, Washimi
As reported by Yamamoto-san ( see klog http://klog.icrr.u-tokyo.ac.jp/osl/?r=14432) F1 GAS output was satured. So Yesterday I offloaded the GAS output by changing the offset on F1 and F2 (see klog http://klog.icrr.u-tokyo.ac.jp/osl/?r=14434)
I took advantage of the suspended mode of KAGRA to investigate some strange response of the geophone installed on the Type A suspension. The test has been done with EX geophone. Basically, I changed the cutoff of the calibration filter from 10 mHz to 3 mHz and I measured the TF= GEO/(LVDT*w). By moving back to 3 mHz the DC cutoff the phase shift is reduced a lot, but the DC noise is growing up. I turned on the ID with blending at 140 mHz: the loop is stable, but the 148 mHz peack got excited (see klog http://klog.icrr.u-tokyo.ac.jp/osl/?r=14443).
- I worked in the measurements of the decay times of SR2. I still have to analyze data. No results to show yet.
Washimi-san reported the feedback signal in SRM F0 was too high. I offloaded onto the fishing rod this morning (klog entry 14436). According to the VIS wilki fishing rod page, this fishing rod does not suffer of any known problems.
- I will postpone asking questions about inertial damping again for next week.
- 3D CAD model of the present IMMT2 suspension is completed.
- Working on the Type-A situation reproduction, which was assigned at cryo-payload upgrade review. (Shoda)
- Comparing the IP performance from the design.
- Test parts have been delivered.
- The experiment will start from the week after next week.
- Test of the LVDT card modification on the grounding is planed.
- KAGRA is still in the suspended mode. Only remote measurements are available until May 31st (Q-measurement by Fabian, Type-A measurement by Lucia).
- More activities in the control room and the site will be available from June 1st (Noise budget by Ushiba).
- Decision for O3c in Virgo will be done within two weeks (from EO). If O3c was decided, task for O3c will be available as well (e.g. Sensor correction for SRs).
- Post Obs. task has priority higher than task for O3c. Two weeks, at least, in June will be assigned for Post Obs. task.
- The site work towards O4 will be rescheduled.
- Mechanics of Type-A tower + Cryo-payload: Shoda (on going)
- Performance of Type-A: Lucia
- Control of Type-A: Lucia?
- Control (Sensor correction): Fujii
- Type-B: Fabian (on going)
- Personnel towards O4
- Takahashi supervises both teams.
- Type-A team: Sato, Miyo, Takahashi, company person
- Type-B/Bp team: Shoda(supervisor), Hirata, Fabian, Terrence, Washimi, company person
- On 2020/6/5(Fri)